OpenStudy (liizzyliizz):

H(z)= 2^(1-6z) find the derivative. i started this but when i checked it on the calculator, i have no idea how it got the answer --12*ln(2)*(1/64)^x

OpenStudy (liizzyliizz):

-12* sorry i didnt mean to put - twice

OpenStudy (anonymous):

let y = 2^(1-6z) ln y = (1-6z) ln 2 dy/dz * 1/y = -6 ln 2 dy/dz = -6 ln 2 * y = -6 ln 2 * 2 ^(1-6z) = - 6 ln 2 *( 2^1 * 2 ^ -6z) = -12 ln2 * (1/64^z)

OpenStudy (liizzyliizz):

is putting the ln automatically something i should of known about? O.o i feel like there are rules that i either dont know or have forgotten, that are making this seem much harder then what it really is.

OpenStudy (anonymous):

you take logs ( ie ln) of each side when the unknown variable is in the exponent - in this case 1- 6z. a simpler example is finder the derivative of 2^x: let y = 2^x then ln y = ln 2^x ln y = x ln 2 differentiate both sides: dy/dx * 1/y = d (x ln 2) / dx = ln 2 dy/dx = ln2 * y = 2^x * ln 2

OpenStudy (anonymous):

d (ln y) / dx = dy/dx * 1/y is by the chain rule - the derivative of ln y = 1/y and derivative of y is dy/dx

OpenStudy (anonymous):

derivative of y with respect to x is dy/dx

OpenStudy (liizzyliizz):

ooooh ok